Find the slope of each line:
1.through (2,5) and (7,5)
2.2x-5y=4
3.through (5,1) and (7,2)
4.x-2=0
5. A line parrallel to 3x-2y=5

Answer by Mathtut(3670)

slope = rise/run or y2-y1/x2-x1 where (x1,y1) and (x2,y2)are points on the line
also lines that are parallel have equal slopes and in order to see the slope in an equation we need to write it in y=mx+b format where m is the slope and b is the y intercept.
1.(5-5)/(7-2)=0 slope =0 which is a horizontal line
2.2x-5y=4 add 5y and subtract 4 to each side 5y=2x-4 divide by 5 y=2/5x-4/5
therefore slope =2/5
3.(2-1)/(7-5)=1/2 ---> slope=1/2
4.x=2 undefined slope . this is a vertical line
5.line parallel to 3x-2y=5 add 2y, subtract 5 and divide by 2 on each side
y=3/2x-5/2 slope is 3/2 and since line is parallel also has a slope of 3/2